Managerial Accounting Benliver, Inc. manufactures and distributes a number or products to retailers. One of these products, Super Jetdry, requires
Answer #1

Answer

a) Sales Budget

July August September Quarter
Sales in Unit 48,000 60,0000 7,5000 1,83,000
Selling Price $47.50 $47.50 $47.50 $47.50
Total Sales $22,80,000 $28,50,000 $35,62,500 $86,92,500

b) Production Budget

July August September Quarter
Sales In unit 48,000 60.000 75.000 1,83,000
Add: Ending Inventory

15,000

18,000 11000 44,000
(3000+20% of 60,000) (3,000+20% of75,000) (3000+20% of 40,000)
Needed 63,000 78,000 86,000 2,27,000
Less: Beginning Inventory 12,600 15,000 18,000 45,600
Unit To be Produced 50,400 63,000 68,000 1,81,400

C) Direct Material Purchase Budget

July Auguest Septumber Quarter
Unit to be produced 50,400 63,000 68,000 1,81,400
Ponds 1,00,800 1,26,000 1,36,000 3,62,000
(50400*2*) (63,000*2) 68,000*2) (1,81,400*2)
Ending Inventory ($0%of next month production) 50,400 54,400 30,400** 1,35,200
Needed 151200 180400 1,66,400 4,97,200
Beginning Inventory 40,320 50,400 54,400 1,45,120
Purchases 110880 130000 112000 351880

Purchase cost

July Auguest September Quarter
Purchase Ponds 110880 120000 112000 351880
Purchase rate per pond $3.5 $3.5 $3.5 $3.5
Purchase Cost $3,88,080 $4,20,000 $3,92,000 $12,31,580

* means every 1 Unit of Finesh goods required 2 Ponds of raw material

calculation production required in July is 50,400 Opening stock required 40 % of Next month production that is 20160 clossing stock of june is 40,320/- means is 40320/20160 = 2 ponds

** Calculation of Opening Stock of Sept.

Calculation of Production of Oct. = sales + clossing Stock - Openig Stock

=40,000+3,000+20%of 30,000( Nov sales)-11000

=38000 units

Raw Material Required = 38000*2 = 76,000

Openig Stock is 30,400/-( 76,000 *40%)